Project

General

Profile

Run Plan » History » Version 751

Version 750 (Jonathan Asaadi, 01/11/2018 12:32 PM) → Version 751/754 (Jason St.john, 01/15/2018 07:20 PM)

{{toc}}

h1. Run Plan (Run IV)/Pixlar

Please help keep this list complete and up to date.



h2. Important notes

* *Always: Please Keep F:MC6IC at or below 7E8 ppp* (for halo management reasons)
%{color:blue} Communicate with the MCR to keep the primary protons at this device close to this level.%

* *%{color:blue} If you are the evening shifter and there is no now owl shifter coming on after you, you please do the following:%* following%*

# Call MCR at the end of your shift and tell them we won't be taking beam overnight
# Tell them there will be no shifter till 8am but we will be running our DAQ overnight taking cosmics data
# If they need to contact someone please call Andrea (1-630-400-4780)
# Start a new run before you leave that will hopefully last the night

+%{color:blue} *Note: Current issues%*+ :

* %{color:red} *Alarms% to be ignored*:
** *CAEN Boards 10 and 11 are not streaming data blocks at the moment*
** *When we have no beam* Magnet and Beam info can be red in Synoptic.



h2. Planned beam interruptions

|_. ongoing/done |_. Time and duration |_. Reason |_. notes |
|*ongoing*| Wednesday 2018-01-10 06:00 - 17:00 | Beam maintenance |open access |
|*Done*| Tuesday 2017-12-19 06:00 - 12:00 | Beam maintenance |open access |
|*Done*| Thursday 2017-12-7 06:00 - TBD | Beam down due to septum problems |open access |
|*Done*| Tuesday 2017-06-27 06:00 - 15:00 | Beam repairs | TBA |
|*Done*| Thursday 2017-05-11 08:00 - 09:00 | Argon Delivery | Open access|
|*Done*| Wednesday 2017-05-15 06:00 - 18:00 | Installation in MC6 | Beam not ensured |
|*Done*| Monday 2017-05-15 06:00 - 18:00 | Work on the accelerator | Control access|
|*Done*| Thursday 2017-05-11 08:00 - 09:00 | Argon Delivery | Open access|
|*done*| Thursday 2017-05-04 07:00 - 16:00 | Installation upstream & argon delivery | Open access|
|*done*| Tuesday 2017-04-25 07:00 - Thursday 2017-04-26 16:00 | accelerator maintenance days & argon delivery| Open access|
|*done*| Thursday 2017-04-20 08:00 - 09:00 | Argon Delivery | Open access|
|*done*| Sunday 2017-04-16 12:00 - Wednesday 2017-04-19 12:00 | Replace the on the B tank for the MI52 septa | Open access|
|*done*| Thursday 2017-04-13 08:00 - 12:00 | Argon Delivery & other stuff | Open access|
|*done*| Thursday 2017-04-06 08:00 - 09:00 | Argon Delivery | Open access|
|*done*| Friday 2017-03-31 06:00 - 18:00 | Work on the accelerator | Open access|
|*done*| Friday 2017-03-24 12:00 - Wednesday 2017-03-30 12:30 | Warm up for filter regeneration| Open access |
|*done*| Thursday 2017-03-24 08:00 - 09:00 | Argon Delivery | Open access|
|*done*| Wednesday 2017-03-22 07:00 - 16:00 | Linac repair | Control access|
|*done*| Tuesday 2017-03-21 07:00 - 16:00 | Work at MCcenter | Open access|
|*done*| Monday 2017-03-20 07:00 - 16:00 | Work at MCcenter | Open access|
|*done*| Thursday 2017-03-16 08:30 - 09:30 | Argon Delivery | Open access|
|*done*| Wednesday 2017-03-15 07:00 - 17:00 | Work at MCcenter | Open access|

h2. In case the beam goes down

This condition is now handled automatically. Do nothing special unless the DAQ will not run. As usual, call an expert if you cannot solve the problem quickly on your own.

Missing Switchyard signals $30 and $36 should be faked by an ACNET script if they are dropped from the timeline:
* If the last SWYD was < 120 seconds ago, go to/stay in Beam On mode
* If the last SWYD was >120 seconds ago, go to/stay in Beam Off Mode by sending fake $30 and $36 based on $00 and delays

h2. Current Runs

*If asked by MCR what kind of beam you'd like please specify these:*

* @MC6IC@ around 7E8 (see table below for run-specific value)
* Steady intensity during the spill: Linear rise on @MC6IC@ during the spill.
* Energy of 80 GeV for proton.
* @MC7U09@ and @MC7U10@ minimized while
* @MC7U11@ and @MC7U12@ maximized

For the other parameters, see the table below.

(If a 'quantity' i.e. BEAMON, is the first in the list, and doesn't have a sign (+ -) in front of it, it's assumed to be positive).
|_. Beam Request |_. Tertiary Beam Magnet Polarity/Current |_. Trigger Configuration |_. Components |_. starting Run |_. starting datetime |_. Notes |
| 64GeV secondary beam, ~7e8 @MC6IC | +100 A |
BEAMON+USTOF+DSTOF+WCCOINC3OF4
ARCLIGHT
FWPULSER+PEDESTALON
COSMICON+COSMIC-BEAMON| V1740 V1740B V1751 V2781 V1495 TDC LARASIC BERN | run 15034 | 2018.01.11 | ARCLIGHT runs *Ongoing* Going back to the first bunch of data we took |
| 64 GeV secondary beam, ~7e8 @MC6IC | -100A |
FWPULSER PEDESTALON
COSMICON+COSMIC-BEAMON
BEAMON+USTOF+DSTOF+WCCOINC3OF4
ARCLIGHT| V1740 V1740B V1751 V2781 V1495 TDC LARASIC BERN | run 14967 | 2018.01.02 19:30PM CST | ARCLIGHT runs *OK* |
| 80 GeV secondary beam, ~7e8 @MC6IC | +20A |
FWPULSER PEDESTALON
COSMICON+COSMIC-BEAMON
BEAMON+USTOF+DSTOF+WCCOINC3OF4
ARCLIGHT| V1740 V1740B V1751 V2781 V1495 TDC LARASIC BERN | run 14754 | 2017.12.24 06:30PM CST | ARCLIGHT runs *OK* (except for 31.12-04.01 - parallel stream data missing) |
| 80GeV secondary beam, ~7e8 @MC6IC | -20A |
BEAMON+USTOF+DSTOF+WCCOINC3OF4
ARCLIGHT
FWPULSER+PEDESTALON
COSMICON+COSMIC-BEAMON| V1740 V1740B V1751 V2781 V1495 TDC LARASIC BERN | run 14648 | 2017.12.15 11:30AM CST | ARCLIGHT runs *OK* |
| 64GeV secondary beam, ~7e8 @MC6IC | +100 A (+85A until magnet cools down) |
BEAMON+USTOF+DSTOF+WCCOINC3OF4
ARCLIGHT
FWPULSER+PEDESTALON
COSMICON+COSMIC-BEAMON| V1740 V1740B V1751 V2781 V1495 TDC LARASIC BERN | run ?? | 2017.12.11 ? | ARCLIGHT runs *OK* |
| 80GeV secondary beam, ~7e8 @MC6IC | -20 A |
FWPULSER+PEDESTALON-SCINTGATE
COSMICON+COSMIC-BEAMON
BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1+PUNCH
COSMICON+LARRY-BEAMON
COSMICON+MICHEL-BEAMON | V1740 V1740B V1751 V2781 V1495 TDC WUT LARASIC | run 13548 | 2017.07.01 | *DONE* Electron-Gamma hunting |
| 80GeV secondary beam, ~7e8 @MC6IC | -40 A |
FWPULSER+PEDESTALON-SCINTGATE
COSMICON+COSMIC-BEAMON
BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1+PUNCH
COSMICON+LARRY-BEAMON
COSMICON+MICHEL-BEAMON | V1740 V1740B V1751 V2781 V1495 TDC WUT LARASIC | run 13404 - run 13547 | 2017.06.23 | *DONE* |
| 80GeV secondary beam, ~7e8 @MC6IC | -100 A |
FWPULSER+PEDESTALON-SCINTGATE
COSMICON+COSMIC-BEAMON
BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1+PUNCH
COSMICON+LARRY-BEAMON | V1740 V1740B V1751 V2781 V1495 TDC WUT LARASIC
Missing: %{color:red}V1742% | run 11572 | 2017.05.16 17:00:00 | *DONE* Antiproton hunting and general beam physics |
| 80GeV secondary beam, ~7e8 @MC6IC | -60 A |
FWPULSER+PEDESTALON-SCINTGATE
COSMICON+COSMIC-BEAMON
BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1+PUNCH
COSMICON+LARRY-BEAMON | V1740 V1740B V1751 V2781 V1495 TDC WUT LARASIC
Missing: %{color:red}V1742% | run 11533 | 2017.05.12 6:30:00 | *DONE* Antiproton hunting and general beam physics |
| 80GeV secondary beam, 1e10 @MC6IC (on/off between runs) | +100A |
FWPULSER-BUSYC
| V1751 V2781 V1495 | run 11522 - run 11532 | 2017.05.11 20:00:00 | *DONE* Special PMT pulser runs |
| 80GeV secondary beam, ~7e8 @MC6IC | -100 A |
FWPULSER+PEDESTALON-SCINTGATE
COSMICON+COSMIC-BEAMON
BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1+PUNCH
COSMICON+LARRY-BEAMON | V1740 V1740B V1751 V2781 V1495 TDC WUT LARASIC
Missing: %{color:red}V1742% | run 11447 | 2017.05.05 13:33:20 | *DONE* Antiproton hunting and general beam physics |
| 80GeV secondary beam, ~7e8 @MC6IC | -100 A |
FWPULSER+PEDESTALON-SCINTGATE
COSMICON+COSMIC-BEAMON
BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1+PUNCH
COSMICON+LARRY-BEAMON | V1740 V1740B V1751 V2781 V1495 TDC WUT LARASIC
V1742 | run 11416 | 2017.05.02 11:00:00 | *DONE* Antiproton hunting and general beam physics |
| 80GeV secondary beam, ~7e8 @MC6IC | -40 A |
FWPULSER+PEDESTALON-SCINTGATE
COSMICON+COSMIC-BEAMON
BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1+PUNCH
COSMICON+LARRY-BEAMON | V1740 V1740B V1751 V2781 V1495 TDC WUT LARASIC
V1742 | run 11383 - run 11415 | 2017.04.28 16:55:00 | *DONE* Electron-Gamma hunting |
| 80GeV secondary beam, ~5e8 @MC6IC | -100 A |
FWPULSER+PEDESTALON-SCINTGATE
COSMICON+COSMIC-BEAMON
BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1+PUNCH
COSMICON+MICHEL-BEAMON
COSMICON+LARRY-BEAMON | V1740 V1740B V1751 V2781 V1495 TDC WUT LARASIC
V1742 | run 11376 - run 11382 | 2017.04.27 16:22:00 | *DONE* Trial antiproton search & general beam physics |
| 80GeV secondary beam, ~5e8 @MC6IC | -40 A |
FWPULSER+PEDESTALON-SCINTGATE
COSMICON+COSMIC-BEAMON
BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1+PUNCH
COSMICON+MICHEL-BEAMON
COSMICON+LARRY-BEAMON | V1740 V1740B V1751 V2781 V1495 TDC WUT LARASIC
V1742 | run 11326 - run 11375 | 2017.04.20 17:35 | *DONE* Electron-Gamma hunting |
| 80GeV secondary beam, ~5e8 @MC6IC | -20 A |
FWPULSER+PEDESTALON-SCINTGATE
COSMICON+COSMIC-BEAMON
BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1+PUNCH
COSMICON+MICHEL-BEAMON
COSMICON+LARRY-BEAMON | V1751 V2781 V1495 TDC WUT LARASIC V1742
Missing: %{color:red}V1740 V1740B% | run 11325 | 2017.04.20 16:22:52 | *DONE* Electron TOF calibration sample |
| 80GeV secondary beam, ~5e8 @MC6IC | -40 A |
FWPULSER+PEDESTALON-SCINTGATE
COSMICON+COSMIC-BEAMON
BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1+PUNCH
COSMICON+MICHEL-BEAMON
COSMICON+LARRY-BEAMON | V1751 V2781 V1495 TDC WUT LARASIC V1742
Missing: %{color:red}V1740 V1740B% | run 11323-11324 | 2017.04.20 15:13:00 | *DONE* Electron TOF calibration sample |
| 80GeV secondary beam, ~5e8 @MC6IC | -40 A |
FWPULSER+PEDESTALON-SCINTGATE
COSMICON+COSMIC-BEAMON
BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1+PUNCH
COSMICON+MICHEL-BEAMON
COSMICON+LARRY-BEAMON | V1740 V1740B V1751 V2781 V1495 TDC WUT LARASIC
V1742 | run 11309 | 2017.04.18 17:45:00 | *DONE* Electron-Gamma hunting |
| 80GeV secondary beam, <2e9 @MC6IC | -40 A |
FWPULSER+PEDESTALON-SCINTGATE
COSMICON+COSMIC-BEAMON
BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1+PUNCH
COSMICON+LARRY-BEAMON | V1740 V1740B V1751 V2781 V1495 TDC WUT LARASIC
V1742 | run 11191 sub-run 431 | 2017.04.11 12:15:51 | *DONE* Electron-Gamma hunting |
| 80GeV secondary beam, <2e9 @MC6IC | -20 A |
FWPULSER+PEDESTALON-SCINTGATE
COSMICON+COSMIC
BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1+PUNCH
COSMICON+LARRY-COSMIC | V1740 V1740B V1751 V2781 V1495 TDC WUT LARASIC
%{color:red}-V1742-% | run 11185 sub-run 153 | 2017.04.10 12:18:51 | *DONE* Electron-Gamma hunting |
| 80GeV secondary beam, <2e9 @MC6IC | +100 A |
FWPULSER+PEDESTALON-SCINTGATE
COSMICON+COSMIC
BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1+PUNCH
COSMICON+LARRY-COSMIC | V1740 V1740B V1751 V2781 V1495 TDC WUT LARASIC
%{color:red}-V1742-% | run 11148 | 2017.04.06 18:54:06 | *Done*
General beam physics |
| 80GeV secondary beam, <2e9 @MC6IC | +100 A |
FWPULSER+PEDESTALON-SCINTGATE
COSMICON+COSMIC
BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1+PUNCH
COSMICON+MICHEL-COSMIC
COSMICON+LARRY-COSMIC | V1740 V1740B V1742 V1751 V2781 V1495 TDC WUT LARASIC | run 10994 | 2017.03.30 10:42:58 | *DONE*
General beam physics |
| | %{color:red}STOP FOR CRYOFILTER REGENERATION% | | | | | |
| None | + 100 A (but doesn't matter) |FWPULSER ~BUSYA BUSYB BUSYC |1742 V1751 V2781 V1495| run 10983 | 2017.03.24 13:32:21 | *DONE*
LAr "Afterglow" runs using the last of the LAr before a stop |
| 80GeV secondary beam, <2e9 @MC6IC | +100 A |
FWPULSER+PEDESTALON-SCINTGATE
COSMICON+COSMIC
BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS1
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS1+PUNCH
COSMICON+MICHEL-COSMIC
COSMICON+LARRY-COSMIC | V1740 V1740B V1742 V1751 V2781 V1495 TDC WUT LARASIC | run 10831 | 2017.03.15 03:35:48 | *DONE*
General beam physics|

h2. TPC Voltages

*FOR 3mm RUN*
Based on e-log entry "&#35;2544":http://dbweb0.fnal.gov/ECL/lariat/E/show?e=2544:
| Plane | Setting | Readback |
| Shield | -298V | -298.75V |
| Induct | -27V | -18.5V |
| Collect | +340V | +338.4V |
| Cathode| - 23,500V| |

*FOR 5mm RUN*
Based on e-log entry "&#35;14547":http://dbweb5.fnal.gov:8080/ECL/lariat/E/show?e=14547:
| Plane | Setting | Readback |
| Shield | -325V | -325.75V |
| Induct | 0V | -0.63V |
| Collect | +423V | +421.88V |
| Cathode| - 23,500V| |

*FOR Pixlar RUN*

| Plane | Setting | Readback |
| Shield | -400V | |
| Induct | 0 V | |
| Collect | 0 V | |
| Cathode| - 23,500V| |

h2. Cryo PMT voltages

Based on e-log entry "&#35;11099":http://dbweb0.fnal.gov/ECL/lariat/E/show?e=11099:
| PMT | Slow ramp setting | Final readback |
| HMM | +1400V | +1382V |
| ETL | +1225V | +1216V |
|Aerogel | -1300V | -1298V|

h1. Investigations and Mysteries

h2. Things to check:

# Visual inspection and documentation of the as-is
# Fix the problem

h2. Checkout before buttoning up
# %{color:red} Before doing any work on TPC, you MUST lock out the Glassman power supply in MCenter control room prior to this test with a RED LOTO lock, done by a LOTO II qualified person.%
# Confirm correct correspondence of bias feedthrough connectors to end of cable, before bias board.
# ✓ Re-insert TPC if needed.
# ✓ Three Lariateers independently confirm firm mechanical connection (and pin alignment!) on all 10 motherboards
# Is there a ground connection for the Last Field Cage Tube? Check it.
# Confirm filter pot inline resistances with blue DVM
# With stop sign in place, ohmic test to confirm field cage is in the circuit again.
# Confirm collection voltage is connected to collection wires
# Measure VDD and VDDP voltages at the front panel of the WRD boards. If they are more than 25 mV off
from 1.8V this indicates a potentially serious problem -- contact Dean for advice. Post values in eLog
# With Bias wire voltages on, measure current flow and verify amps match expected values
# Take a teensy channel-by-channel pulser run and verify all channels are firing where expected
# Turn on the teensy pulser and page through all 480 channels with the waveform program. Check for
abnormal gains and shapes
# Check for ground shorts between cathode cable and ground. Glassman power
supply in MCenter control room should still be locked by red LOTO lock.
# Throughout the close up checkout process, take several pedestal runs and look for dead and noisy
channels

h2. Other investigations

* FIXED: *2017.03.23* Why did the induction plane draw 80 microamps at ~ -18 V, when the collection and shield, at a few hundred volts negative and positive, draw ~microamps? Because we mounted the induction in the wrong way!

h1. [[Old Run Plan from Run II]]